OpenStudy (anonymous):

Find the equation of the circle with center at (-4,6) and passes through point (-4,7)

OpenStudy (therealmeeeee):

equation of a circle in a center (h , k ) / radius r form is: (x - h)^2 + (y - k)^2 = r^2 center: (-5 , 6) (x + 5)^2 + (y - 6)^2 = r^2 if it passes through the point x = 1 and y = 3 (1 + 5)^2 + (3 - 6)^2 = r^2 (6)^2 + (-3)^2 = r^2 36 + 9 = r^2 r = sqrt(45) hence the equation is: (x + 5)^2 + (y - 6)^2 = 45

OpenStudy (therealmeeeee):

is that helping u???????????

OpenStudy (therealmeeeee):

ty